OpenStudy (anonymous):

can you show how to solve 7-9r-(r+12)≤25

OpenStudy (anonymous):

7-9r-(r+12)≤25 =>\[-10r-5\le25\]\[-10r \le30\]\[10r \ge 30\]\[r \ge 3\]

OpenStudy (anonymous):

\[7-9r-(r+12)\le25\] \[7-9r-r-12\le25\] \[7-10r-12\le25\] \[-10r-5\le25\] +5 +5 \[-10r \le30\] /-10 /-10 \[r \ge-3\]

OpenStudy (anonymous):

thanks what about 2x+3<7 after applying the addition property of inequality